OpenStudy (anonymous):

There are x hats on a hat rack. If one hat were to be selected at random, the probability that the hat would be green is . The probability that the hat would be black is . If the only other color remaining is red, what is the total number of red hats in terms of x?

OpenStudy (callisto):

you've missed all the numbers in your question...

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

There are x hats on a hat rack. If one hat were to be selected at random, the probability that the hat would be green is 2/5 . The probability that the hat would be black is 1/5 . If the only other color remaining is red, what is the total number of red hats in terms of x?

OpenStudy (callisto):

no of green hat 2 ------------- = -- x 5 no of green hat = 2x/5 no of black hat 1 ------------- = -- x 5 no of black hat = x/5 no of red hat= total no of hats - no of black hats - no of green hats = x - x/5 - 2x/5 = (5x - x-2x ) /5 = 2x/5

OpenStudy (perl):

@eliassaab can i ask the question here ?